Sally Rogers has decided to invest her wealth equally across the following three​ assets: What are her expected returns and the risk from her investment in the three​ assets? How do they compare with investing in asset M​ alone?   Hint​: Find the standard deviations of asset M and of the portfolio equally invested in assets​ M, N, and O. What is the expected return of investing equally in all three assets​ M, N, and​ O? (round to two decimal places)

States   Probability   Asset M Return   Asset N Return   Asset O Return
Boom   33%   10%   19%   2%
Normal   54%   8%   12%   8%
Recession   13%   2%   -1%   10%
Answer #1

1.Find the expected return of the equally weighted portfolio in the three economic states.

Return of Portfolio in Boom =1/3 (10%) +1/3 (19%) +1/3 (2%) =10.33%

Return of Portfolio in Normal =1/3 (8%) +1/3 (12%) +1/3 (8%) =9.33%

Return of Portfolio in Recession =1/3 (2%) +1/3 (-1%) +1/3 (10%) =3.67%

2. Find the expected returns of Asset M and the Portfolio.

Expected Return Asset M =0.33 ×(10%) +0.54 ×(8%) +0.13 (2%)

E(rM) =3.3% +4.32% +0.26% = 7.88%

Expected Return Portfolio =0.33 ×(10.33%) +0.54 ×(9.33%) +0.13 (3.67%)

E(rp) =3.4% +5.03% +0.47% = 8.498%

3.Find the standard deviation of Asset M and the Portfolio.

Standard Deviation of Asset M =[0.33 ×(0.10 -0.0788)2+0.54 ×(0.08 -0.0788)2+0.13 ×(0.02 -0.0788)2]1/2

=[(0.33 ×0.00044944) +(.54× 0.00000144)+(0.13 ×0.00345744)]1/2

=(0.00059856)1/2=0.00029928 or 2.99%

Standard Deviation of Portfolio =[0.33 ×(0.1033 -0.0849)2+0.54 ×(0.0933 -0.0849)2+0.13 ×(0.367 -0.0849)2]1/2

=[(0.33 ×0.00033856) +(0.54 ×0.00007056) +(0.13 ×0.07958041)]1/2

=(0.0104952805)1/2 = 0.00524764025 or 5.24%

The benefit of the portfolio over Asset M alone is an increase in return of 0.38% and a simultaneous reduction in total risk of 2.25%.